OpenStudy (anonymous):

Solve the equation for w. P = 2l + 2w

OpenStudy (anonymous):

w=(P-2l)/2

jimthompson5910 (jim_thompson5910):

\[\large P = 2l + 2w\] \[\large P - 2l = 2w\] \[\large \frac{P - 2l}{2}=w\] \[\large w=\frac{P - 2l}{2}\]

OpenStudy (anonymous):

Thank you!! :]

OpenStudy (nilankshi):

good